A company is going to invest in a new machine with a capital cost of $250,000. The life of the machine will be 7 years with no salvage value. The machine can produce up to 100,000 parts each of which will make a profit of $2. The operating costs are $25,000 per year. Assuming a MARR of 10%, what is the minimum production level the machine must achieve to be profitable?

between 39,000 and 42,000

greater than 42,000

less than 37,000

between 37,000 and 39,000
Answer #1

Let breakeven be n units, then

EUAC = 250000*(A/P,10%,7) + 25000

= 250000*0.205405 + 25000

= 76351.25

EUAB = 2*n

For breakeven, EUAB = EUAC

2*n = 76351.25

n = 76351.25 / 2 = 38175.62 ~ 38176 (between 37,000 and 39,000)

Last option is correct answer